Find the distance between the points (2, -3) and (5, -4).
Find the distance between the points (-1, 3) and (5, 0). (Hint: Remember to simplify the square root if you can. For example: sqrt 28 = 2 sqrt 7.)
Write your answer to problems 3 and 4 as ordered pairs (x, y).
Find the midpoint between the points (3, 1) and (1, 5).
Find the midpoint between the points (5, 2) and (-1, -2).

Answer by stanbon(75887)

Find the distance between the points (2, -3) and (5, -4).
d= sqrt[(-4-2)^2 + (5-2)^2] = sqrt[36+9] = sqrt(45 = 3sqrt(5)
==============================================================
Find the distance between the points (-1, 3) and (5, 0). (Hint: Remember to simplify the square root if you can. For example: sqrt 28 = 2 sqrt 7.)
Write your answer to problems 3 and 4 as ordered pairs (x, y).
Find the midpoint between the points (3, 1) and (1, 5).
Let the midpoint be (x,y)
x = (3+1)/2 = 2
y = (1+5)/2 = 3
mid-point: (2,3)

Find the midpoint between the points (5, 2) and (-1, -2).
==============
Cheers,
Stan H.
======================